OpenStudy (anonymous):

find the slope of the curve, the line that is tangent to the curve, or the line normal to the curve. 2x^2y -pi(cosy)=3(pi), tangent at (1, (pi))

OpenStudy (anonymous):

so would you do implicit differentiation first?

OpenStudy (anonymous):

yeah for sure

OpenStudy (anonymous):

would it be 4xy(y1)

OpenStudy (anonymous):

\[ 2x^2y -\pi(\cos(y))=3\pi\]right?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

do you know how to do implicit diff?

OpenStudy (anonymous):

i think so... wouldn't it be 4xy(y1) +pi(siny)(y1) ?

OpenStudy (anonymous):

oh no you need the product rule and the chain rule

OpenStudy (anonymous):

think of \[2x^2y -\pi(\cos(y))=3\pi\]as if it was \[2x^2f(x) -\pi(\cos(f(x)))=3\pi\]

OpenStudy (anonymous):

for the first term you need the product rule \[4xf(x)+2x^2f'(x)\]

OpenStudy (anonymous):

for the second one you need the chain rule \[-\pi \cos(f(x))f'(x)\]

OpenStudy (anonymous):

written with \(y\) instead of \(f(x)\) it is \[4xy+2x^2y'+\pi \sin(y)y'=0\]

OpenStudy (anonymous):

ohhhhhh

OpenStudy (anonymous):

so now i just plug in (1,pi)?

OpenStudy (anonymous):

i get the implicit part now! duh. I was so confused O.o

OpenStudy (anonymous):

you can plug in \((1,\pi)\) now and then solve for \(y'\) or you can use algebra to solve for \(y'\) first then plug in the numbers probably easier to do the first method

OpenStudy (anonymous):

good!

OpenStudy (anonymous):

ok so 4(1)(pi) + 2(1)^2(y1) + (pi)sin(1)(y1)=0 4(pi) +4(y1) +

OpenStudy (anonymous):

oh jeez i'm stuck so (pi)sin(1)(y1) would be (pi)(pi/2)(y1)?

OpenStudy (anonymous):

\[4xy+2x^2y'+\pi \sin(y)y'=0\] at \((1,\pi)\) \[4\pi+2y'+\pi\sin(\pi)y'=0\]

OpenStudy (anonymous):

since fortunately \(\sin(\pi)=0\) this is just \[4\pi+2y'=0\] and that is easy to solve for \(y'\)

OpenStudy (anonymous):

ohh i messed up and did sin(1) whoops

OpenStudy (anonymous):

yeah you don't know that one for sure

OpenStudy (anonymous):

y1 =-2pi?

OpenStudy (anonymous):

that is what i get, yes

OpenStudy (anonymous):

yaaas !! thanks so much

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

urmm are u busy cuz i have 1 more quick question?

OpenStudy (anonymous):

sure go ahead and ask

OpenStudy (anonymous):

ok this question is "If the fuction is not differentiable at the given value of x, tell whether the problem is a corner, cusp, vertical tanget or discontinuity" y= -7 -\[\sqrt[3]{x}\] at x=0 so wouldn't I just take the derivitive which is 1/3(x^3) ? and plug in 0 ?

OpenStudy (anonymous):

it's y= -7 - 3sqrtx

OpenStudy (anonymous):

your derivative is wrong

OpenStudy (anonymous):

*slams head on desk*

OpenStudy (anonymous):

\[y=7-\sqrt[3]{x}\] \[y'=-\frac{1}{3\sqrt[3]{x^2}}\]

OpenStudy (anonymous):